clue 1: i am a two digit number
clue 2: my second digit is the square of the first
clue 3: the prodcut of my digits is a cube

Found 2 solutions by jorel1380, Edwin McCravy:
Answer by jorel1380(3719)

24
2^2=4
2x4=8=2^3
The number is 24..

Answer by Edwin McCravy(20081)

clue 1: i am a two digit number
clue 2: my second digit is the square of the first
clue 3: the product of my digits is a cube
Looking at clue 2:

The only digits whose squares are also digits are 0,1,2, and 3 and their
squares are  0,1,4, and 9, respectively.  So the only possibilities are 
00, 11, 24, and 39.   

Looking at clue 1:

We rule out 00 as being a two-digit number.

So the only candidates are 11, 24, and 39.  

Looking at clue 3:

They must be so that their product is a cube. 

It's true of 11 because 1×1 = 1 = 1³
It's true of 24 because 2×4 = 8 = 2³
It's true of 39 because 3×9 = 27 = 3³

So there are three possible different answers 11, 24, and 39.

Edwin